# New insights into the distribution of the topmost gap in random walks and Lévy flights

### v1

@@ Abstract
Random walks, Lévy flights, Ladder variables, Wiener-Hopf
@@
In [1], we derived quantitative estimates for the asymptotic tail behaviour of $f_{H}(x)$, which led us to classify step distributions into three distinct categories: superexponential, exponential, and subexponential. We also performed a detailed analysis of the case where the step distribution follows a stable law and computed the moments of $H$ for a Gaussian walk. Moreover, we obtained explicit expressions for $f_{H}(x)$ in specific cases, notably for the family of symmetric Erlang distributions.
Here, we establish related results concerning the gap $Z$ between the largest and second-largest positions of a formally infinitely long walk. Our approach bypasses the more intricate computations required in earlier studies [2, 3], which aimed to derive the joint distribution of $Z$ and the time interval between the largest and second-largest positions. The notations of [1] will be used throughout.
Our main results are as follows:
@@
    An interpretation of the topmost gap $Z$ as the smaller of two independent copies of $H$, the first positive position of the walk, leading to a direct derivation of an expression for the gap density $f_{Z}(x)$ in terms of $f_{H}(x)$.
@@
    A systematic characterisation of the asymptotic tail behaviour of $f_{Z}(x)$, for the three different classes of step distributions.
@@
The topmost gap $Z$, defined as the distance between the largest and second-largest positions of a very long walk, is a simple example of extreme order statistics (see [4] for an overview and [5] for a recent probabilistic approach). The key observation of our approach is that $Z$ can be viewed as the smaller of two independent copies of $H$, the first positive position of the walk:
@@
This interpretation is illustrated in figure 1. In each panel, the black point represents the absolute maximum, i.e., the highest point of a formally infinite walk. The distance $H_{1}$ corresponds to the first positive position of the walk when read from left to right, starting from the red point. In this case, this position is reached after two steps. Similarly, $H_{2}$ represents the first positive position when the walk is read from right to left, starting from the green point. Here, this position is reached after a single step. In the first panel, where $H_{1}<H_{2}$, the red point is the second-highest, so $Z=H_{1}$. In the second panel, where $H_{2}<H_{1}$, the green point is the second-highest, giving $Z=H_{2}$. [^1] Notice that the red point is a local maximum of the walk, whereas the green one is not. In general, the second-largest position is a local maximum when it is separated from the largest position by at least two steps, which occurs with probability $1/2$ [1]. To the best of our knowledge, the relation (3) was not previously known in the fluctuation theory of random walks.
@@
Asymptotic tail behaviour
@@
    For Lévy flights ($20<\theta<$ 2), we have [1]
@@
Integrating over $r$ amounts to adding up the residues of the poles at $r=0$ and $r=-p$. We thus obtain
@@
Integrating over $r$ still amounts to summing the residues of the poles at $r=0$ and $r=-p$. The latter is now a double pole, whose residue involves a derivative with respect to $p$, which we denote by an accent. This gives
@@
For step distributions such that $32<\theta<$ 3, $\langle Z^{2}\rangle$ is finite, although $\ell$ is divergent. Obtaining an integral expression for $\langle Z^{2}\rangle$ in that case requires greater care. We shall not elaborate on this.
Higher moments and Laplace transform
@@
Integrating over $r$ amounts to adding up the residues of the poles at $r=0$ and $r=-(p+s)$. We thus obtain
@@
Symmetric Erlang distributions
For this family of step distributions, the step density is given by
@@
The family (61) provides an interesting framework where a wide range of explicit results can be derived, following the approach of [1]. These distributions indeed belong to the class of distributions whose Laplace transform $\hat{\rho}(p)$ is a rational function [^2] As mentioned in [1], the observation that Wiener-Hopf integral equations such as (12) associated with such distributions are solvable by elementary means dates back at least to the works of Wick [14] and Chandrasekhar [15].. The product formula [^3] Here and throughout the following, sums and products over the zeros $z_{b}$ run over the range $b=1,\dots,M-1$, which is empty for $M=1$.
@@
plays a key role in the following. In addition to a double zero at the origin, $\phi(p)$ has $M-1$ zeros with positive real parts, denoted as $z_{b}$ for $b=1,\dots,M-1$, along with their opposites. These zeros sit near the circles with unit radii centered at $p=\pm 1$ (see Figure 2). They have been investigated in detail in [16]. We have in particular
@@
The density of the gap $Z$ is thus the product of the decaying exponential $2{\rm e}^{-2x}$ by a polynomial of degree $2(M-1)$. In (67) and (68), the $S_{k}$ are the elementary symmetric functions of the variables $z_{b}-1$, defined by the identity
@@
The moments of the gap distribution are investigated in the Appendix. Closed-form expressions for $\langle Z^{2}\rangle$ and $\langle Z^{3}\rangle$ are given in (99) and (Third moment).
@@
Appendix: Gap moments for Erlang distributions
@@
along a vertical contour with $10<\mathop{{\rm Re}}u<$ 1, the sums over $k$ and $l$ can be performed separately by using (69), yielding
@@
The general structure of the gap moments appears clearly from (99) and (Third moment). The expression of $\langle Z^{n}\rangle$ involves, in addition to an explicit polynomial in the Erlang parameter $M$ and the extrapolation length $\ell$, a simple sum over the zeros $z_{b}$, a double sum, and so on, up to a sum over $(n-1)$-uples of different zeros.
Large-$M$ asymptotics
@@
This leading-order estimate is however not accurate enough in most practical cases. For a large but finite $M$, the positions of the zeros near the origin, i.e., for $b\ll M$ or $M-b\ll M$, deviate significantly from the uniform distribution mentioned above. To be specific, let us consider the range $b\ll M$. Retaining the notations of [16, Sec. 6.2], and defining

### v2

@@ Abstract
Random walks, Lévy flights, Ladder variables, Wiener-Hopf method
1 Introduction
Extreme-value theory and order statistics have been the subject of extensive study in probability theory [1, 2, 3, 4, 5], statistical physics [6, 7, 8, 9, 10], astrophysics [11], environmental sciences [12], finance [13, 14], reliability theory [15], and other fields [16, 17]. Classical results mostly concern independent and identically distributed random variables [18, 2, 19].
The study of order statistics for one-dimensional random walks was initiated by Pollaczek [20], Wendel [21], and others. More broadly, the fluctuation theory of random walks, developed in the 1950s and 1960s by Sparre Andersen, Spitzer, Baxter, Feller, and Wendel, describes the distributions of extrema and the random times at which they are first or last attained.
Recent work [22] on random walk order statistics has focused on the spacings between the successive ordered positions of the walker. For symmetric walks with continuous increments, the authors derive new limit laws and asymptotic expressions. In particular, when the step distribution is symmetric with finite variance, the expected spacings converge at long times to explicit integral formulas involving the Fourier transform of the step density. This question is further explored in [23], which also examines the long-time behavior of these spacings [^1] We refer to the latter reference for more details on the historical aspects of random walk extreme order statistics. The beginning of the present introduction borrows from it..
Here, we establish new results concerning the topmost gap $Z$, defined as the distance between the largest and second-largest positions of a formally infinitely long walk. Our approach bypasses the more intricate computations required in earlier studies [24, 25], which aimed to derive the joint distribution of $Z$ and the time interval between the largest and second-largest positions.
@@ 1 Introduction
In [26], we derived quantitative estimates for the asymptotic tail behaviour of $f_{H}(x)$, which led us to classify step distributions into three distinct categories: superexponential, exponential, and subexponential. We also performed a detailed analysis of the case where the step distribution follows a stable law and computed the moments of $H$ for a Gaussian walk. Moreover, we obtained explicit expressions for $f_{H}(x)$ in specific cases, notably for the family of symmetric Erlang distributions. The notations of [26] will be used throughout.
Our main results concerning the topmost gap $Z$ between the largest and second-largest positions of a very long random walk are as follows:
@@
    An interpretation of $Z$ as the smaller of two independent copies of $H$, the first positive position of the walk, leading to a direct derivation of an expression for the gap density $f_{Z}(x)$ in terms of $f_{H}(x)$.
@@
    A systematic characterisation of the asymptotic tail behaviour of $f_{Z}(x)$, for the three different classes of step distributions mentioned above.
@@ 2 Distribution of the gap
The key observation of our approach is that $Z$ can be viewed as the smaller of two independent copies of $H$, the first positive position of the walk:
@@
This interpretation is illustrated in figure 1. In each panel, the black point represents the absolute maximum, i.e., the highest point of a formally infinite walk. The distance $H_{1}$ corresponds to the first positive position of the walk when read from left to right, starting from the red point. In this case, this position is reached after two steps. Similarly, $H_{2}$ represents the first positive position when the walk is read from right to left, starting from the green point. Here, this position is reached after a single step. In the first panel, where $H_{1}<H_{2}$, the red point is the second-highest, so $Z=H_{1}$. In the second panel, where $H_{2}<H_{1}$, the green point is the second-highest, giving $Z=H_{2}$. [^2] Notice that the red point is a local maximum of the walk, whereas the green one is not. In general, the second-largest position is a local maximum when it is separated from the largest position by at least two steps. This event occurs with probability $1/2$ [26]. To the best of our knowledge, the relation (3) was not previously known in the fluctuation theory of random walks.
@@
3 Asymptotic tail behaviour
@@ 3.3 Third class
    For Lévy flights ($0<\theta<2$), we have [26]
@@ 4.1 First moment
Integrating over $r$ amounts to adding up the residues of the poles at $r=0$ and r=pr=-p. We thus obtain
@@ 4.2 Second moment
Integrating over $r$ still amounts to summing the residues of the poles at $r=0$ and r=pr=-p. The latter is now a double pole, whose residue involves a derivative with respect to $p$, which we denote by an accent. This gives
@@
For step distributions such that $2<\theta<3$, $\langle Z^{2}\rangle$ is finite, although $\ell$ is divergent. Obtaining an integral expression for $\langle Z^{2}\rangle$ in that case requires greater care. We shall not elaborate on this.
4.3 Higher moments and Laplace transform
@@
Integrating over $r$ amounts to adding up the residues of the poles at $r=0$ and r=(p+s)r=-(p+s). We thus obtain
@@
5 Symmetric Erlang distributions
For this family of distributions, the step density is given by
@@ 5 Symmetric Erlang distributions
The family (61) provides an interesting framework where a wide range of explicit results can be derived, following the approach of [26]. Another recent work [35] is devoted to the statistics of gaps in random walks with the same family of step distributions. These distributions belong to the class of distributions whose Laplace transform $\hat{\rho}(p)$ is a rational function [^3] As mentioned in [26], the observation that Wiener-Hopf integral equations such as (12) associated with such distributions are solvable by elementary means dates back at least to the works of Wick [36] and Chandrasekhar [37].. The product formula [^4] Here and throughout the following, sums and products over the zeros $z_{b}$ run over the range $b=1,\dots,M-1$, which is empty for $M=1$.
@@
plays a key role in the following. In addition to a double zero at the origin, $\phi(p)$ has M1M-1 zeros with positive real parts, denoted as $z_{b}$ for $b=1,\dots,M-1$, along with their opposites. These zeros sit near the circles with unit radii centered at $p=\pm 1$ (see Figure 2). They have been investigated in detail in [38]. We have in particular
@@
The density of the gap $Z$ is thus the product of the decaying exponential $2{\rm e}^{-2x}$ by a polynomial of degree 2(M1)2(M-1). In (67) and (68), the $S_{k}$ are the elementary symmetric functions of the variables $z_{b}-1$, defined by the identity
@@
The moments of the gap distribution are investigated in the Appendix. Closed-form expressions for $\langle Z^{2}\rangle$ and $\langle Z^{3}\rangle$ are given in (99) and (A.3).
@@
Appendix A Gap moments for Erlang distributions
@@ A.1 First moment
along a vertical contour with $0<\mathop{{\rm Re}}u<1$, the sums over $k$ and $l$ can be performed separately by using (69), yielding
@@ A.3 Third moment
The general structure of the gap moments appears clearly from (99) and (A.3). The expression of $\langle Z^{n}\rangle$ involves, in addition to an explicit polynomial in the Erlang parameter $M$ and the extrapolation length $\ell$, a simple sum over the zeros $z_{b}$, a double sum, and so on, up to a sum over (n1)(n-1)-uples of different zeros.
A.4 Large-$M$ asymptotics
@@
This leading-order estimate is however not accurate enough in most practical cases. For a large but finite $M$, the positions of the zeros near the origin, i.e., for $b\ll M$ or $M-b\ll M$, deviate significantly from the uniform distribution mentioned above. To be specific, let us consider the range $b\ll M$. Retaining the notations of [38, Sec. 6.2], and defining